OpenStudy (anonymous):

Factor : 64x^2+16x+1=0

OpenStudy (anonymous):

(8x+1)(8x+1)=0 (8x+1)^2=0

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

you're welcome